a car travel the first 20km witha speed of 4km/hr and the next 40km/hr with a speed of 30km/hr find over all averege
speed
1
Expert's answer
2017-06-13T09:49:09-0400
By the definition, the average speed is the total distance traveled divided by the total time:
v_avg=d_tot/t_tot .
It is obvious that the total distance is equal to
d_tot=d_1+d_2=20 km+40 km=60 km.
Let’s first find the time that the car needs to travel 20 km at 40 km⁄h:
t_1=d_1/v_1 =(20 km)/(40 km/h)=0.5 h.
Similarly, we can find the time that the car needs to travel the next distance of 40 km at 30 km⁄h:
t_2=d_2/v_2 =(40 km)/(30 km/h)=1.33 h.
Then, we can find the total time for the entire journey:
t_tot=t_1+t_2=0.5 h+1.33 h=1.83 h.
Finally, we can find the average speed of the bus for the entire journey:
v_avg=d_tot/t_tot =(60 km)/(1.83 h)=33 km/h.

Answer: v_avg=33 km/h.